The heat of vaporization of benzene (C6H6) is 30.77 kJ/mol at it's boiling point (80 C). As a liquid, it has a heat capacity of 134.8 J/mol*K, and as a gas it has a heat capacity of 82.4 J/mol*K. Calculate the enthalpy change of heating 245g of benzene from 50 C to 100 C.
Answer #1

moles of benzene = 245 / 78 = 3.141

50oC --------------------> 80oC ----------------------> 100oC

               1                         2                 3

Q1 = m C dT

Q1 = 3.141 x 134.8 x (80-50) = 12702 J = 12.70 kJ

Q2 = 30.77 x 3.141   = 96.65 kJ

Q3 = m C dT

Q3 = 3.141 x 82.4 x (100 -80)

Q3 = 5176 J = 5.176 kJ

Q = Q1 + Q2 + Q3

Q = 12.70 + 96.55 + 5.176

Q = 114 kJ

enthalpy change = 114 kJ